# The Sounds of Home: A Speech-Removed Residential Audio Dataset for Sound Event Detection

### v1

@@ 1 Introduction
We release a large mutli-recording audio dataset which is recorded at multiple location within domestic or home environment of older adults. This dataset includes 1342 hours of audio recordings and is intended for sound event detection research.
@@ 2 Dataset Collection Methodology
An objective of this study was to collect a real-world audio dataset from domestic environments to support research on sound event detection in everyday conditions. To achieve this, we conducted a study in collaboration with the Living and Care Lab (LiCalab) at Thomas More University in Geel/Turnhout, Belgium, funded by the VITALISE Transnational Access programme [5].
@@ 2.3 Data Collection
To document the acoustic characteristics of the environments, we created detailed floor plans of each home, including precise measurements. Information about the construction materials used in the homes was also recorded, with the goal of being able to reproduce the acoustic properties of the spaces in future research. After installation, photographs were taken of the recording devices in their respective placements for documentation purposes. the devices were installed by us, with assistance from LiCalab staff [5], to ensure proper setup and minimize potential disturbances to the participants’ daily routines.
@@ 3.3 Threshold Analysis
The audio dataset analysis focused on the distribution of predefined vocal-related labels to determine an optimal confidence threshold that could effectively differentiate spoken voice from ambient sounds while preserving as much of the dataset as possible. The goal was to remove all audio segments containing speech-related labels without significantly reducing the total hours of audio data. To cover all situations in which the audio could contain spoken voice potentially identifiable as personal information, all segments above the threshold were removed that included the labels: “Speech”, “Singing”, “Male singing”, “Female singing”, “Child singing”, “Male speech, man speaking”, “Female speech, woman speaking”, “Conversation”, “Narration, monologue”, “Music”. Figure 2 shows the relationship between various confidence thresholds and the proportion of the dataset containing the predefined vocal-related labels with confidence values exceeding these thresholds.
@@ 3.4 Speech Removal Process
A subsequent cleaning phase involves reading the JSON files and processing the audio files to remove the detected speech segments. The probability threshold of 0.2 is applied to the confidence values of the estimates to determine whether a segment contains speech. The script calculates the start and end times for each detected speech segment, extending these intervals by one second before and after the detected speech to ensure complete removal.
The removal of speech consists of modifying the sections of the audio file identified as containing speech by replacing the identified speech segments with low-amplitude noise, generated with an amplitude of 1e-10. This step is added to prevent division-by-zero errors during subsequent processing and to maintain the structural integrity of the audio file. The modified audio file, now free of speech, is saved to a new location in the temporary storage cluster server. Figure 3 shows an example of the first part of this process applied to a 10-second audio fragment containing spoken voice.
To ensure thorough removal of speech, a two-step process is employed using the two models in cascade. Initially, the audio file is processed, generating an intermediate file allocated in the cluster server. This intermediate file undergoes a second round of processing with a different model. The intermediate file is then deleted to maintain a clean workspace.

### v2

@@ 1 Introduction
We release a large multi-recording audio dataset recorded at multiple residential locations of older adults. This dataset includes 1342 hours of audio recordings and is intended for sound event detection research111The complete dataset is downloadable from the CVSSP server: https://www.cvssp.org/data/ai4s/sounds_of_home. For segmented downloads, the dataset is also available in four parts on Zenodo. The first part is at https://doi.org/10.5281/zenodo.12737915, where links to the remaining parts are provided..
@@ 2 Dataset Collection Methodology
An objective of this study was to collect a real-world audio dataset from domestic environments to support research on sound event detection in everyday conditions. To achieve this, we collaborated with the Living and Care Lab (LiCalab) at Thomas More University in Geel/Turnhout, Belgium, funded by the VITALISE Transnational Access programme [5].
@@ 2.3 Data Collection
To document the acoustic characteristics of the environments, we created detailed floor plans of each home, including precise measurements. Information about the construction materials used in the homes was also recorded, with the goal of reproducing the acoustic properties of the spaces in future research. After installation, photographs were taken of the recording devices in their respective placements for documentation purposes. The devices were installed by us, with assistance from LiCalab staff [5], to ensure proper setup and minimize potential disturbances to the participants’ daily routines.
@@ 3.3 Threshold Analysis
The audio dataset analysis focused on the distribution of predefined vocal-related labels to determine an optimal confidence threshold that could effectively differentiate spoken voice from ambient sounds while preserving as much of the dataset as possible. The goal was to remove all audio segments containing speech-related labels without significantly reducing the total hours of audio data. To cover all situations in which the audio could contain spoken voice potentially identifiable as personal information, all segments above the threshold were removed that included the labels: “Speech”, “Singing”, “Male singing”, “Female singing”, “Child singing”, “Male speech, man speaking”, “Female speech, woman speaking”, “Conversation”, “Narration, monologue”, “Music”. Figure 2 shows the
relationship between various confidence thresholds and the proportion of the dataset containing the predefined vocal-related labels with confidence values exceeding these thresholds.
@@ 3.4 Speech Removal Process
A subsequent cleaning phase involves reading the JSON files and processing the audio files to remove the detected speech segments. A 0.2 probability threshold is applied to confidence values to determine speech presence. The script calculates the start and end times for each detected speech segment, extending these intervals by one second before and after the detected speech for thorough removal.
The removal of speech modifies sections of the audio file containing speech by replacing these with low-amplitude noise, generated at relative amplitude. This step prevents division-by-zero errors during subsequent processing and maintains the audio file’s structural integrity. The modified file, now free of speech, is saved to a new location in the temporary storage cluster server. Figure 3 illustrates this process on a 10-second audio fragment with spoken voice.
To ensure thorough removal of speech, a two-step process is used with two models in cascade. Initially, the audio file is
processed, creating an intermediate file on the cluster server. This file undergoes a second processing with a different model, then is deleted to maintain a clean workspace.